OpenStudy (anonymous):

how do you write 27 ^-2/3 in radical notation

OpenStudy (anonymous):

can anyone help???

OpenStudy (anonymous):

please

OpenStudy (anonymous):

you are working with index notation -2/3 - means that the term is under 1 the top part of the fraction (numerator) is the power Denominator is the root 1/27 cube root of 27 = 3 now 1/3/^2 simplify = 1/9